What is the sum of 1 up to 100
Answer by stanbon(75887)

What is the sum of 1 up to 100
-----
List the numbers: 1,2,3,4,....98,99,100
----
Notice that
1+100 = 101
2+99 = 101
3+97 = 101
........
50+51 = 101
================
Sum = 50(101) = 5050
========================
Cheers,
Stan H.